OpenStudy (anonymous):

Please can someone explain the following (x+4)(x=6)

myininaya (myininaya):

i can't it doesn't make sense to me sorry

OpenStudy (anonymous):

can be written as (x+4)x=6 x^2+4x=6 x^2+4x-6=0 and then can be solved using QE

OpenStudy (anonymous):

sorry if it's not the right answer